OpenStudy (steve816):

I just took a math test, but there was a rigged question. Do not worry everyone, I'm not cheating, I already took the test and asking the question after taking it. It is about trigonometry and I was getting headache attempting the problem.

OpenStudy (steve816):

\[\tan(3x) + 1 = \sec(3x) \]

OpenStudy (steve816):

Find all solutions.

zepdrix (zepdrix):

Rigged? :) lol

OpenStudy (steve816):

I knew that I had to substitute the 3x with just a single x, but I just couldn't solve from there.

zepdrix (zepdrix):

Here is my approach, square both sides,\[\large\rm \tan^2(3x)+2\tan(3x)+1=\sec^2(3x)\]Use your Pythagorean Identity on the right side of the equation,\[\large\rm \tan^2(3x)+2\tan(3x)+1=\tan^2(3x)+1\]And then you have a lot of nice cancellations from there :) Any confusion?

OpenStudy (steve816):

Wait, where did you get the 2tan(3x) And now that you say pythagorean identity, I feel so dumb that I didn't recognize it :(

OpenStudy (steve816):

I'm a failure at math! Nooooooo, my test score is ruined :(

zepdrix (zepdrix):

\[\large\rm (a+1)^2=a^2+\color{orangered}{2a}+1^2\]

zepdrix (zepdrix):

\[\large\rm (\tan3x+1)^2=\tan^2(3x)+\color{orangered}{2\tan3x}+1^2\]

OpenStudy (unklerhaukus):

tan + 1 = sec sin/cos + cos/cos = 1/cos sin + cos = 1 (cos ≠ 0)

OpenStudy (steve816):

ohhhh, I understand now.

zepdrix (zepdrix):

ya sines and cosines is another good approach :)

OpenStudy (steve816):

Alright, thanks guys.